# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 2 Series and generating functions in $\mathbb{C}$
It is the ray in $\mathbb{C}$ going from the origin in the direction $\xi$. Clearly, $\xi\mathbb{R}_{\geq 0}$ is closed under addition. If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$, then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$. If
@@
is a light weight and $n\in\mathbb{N}_{0}$ then $h(W(n))\in\xi^{n}\mathbb{R}_{\geq 0}$. The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties. For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
@@
    absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
@@
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of the well known Hadamard’s formula for the radius of convergence of a formal power series.
@@
We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems, which we state only for the radius of convergence $R\geq 1$. The first theorem is classical, see [1, 2] and [29, Chapter II.7].
@@
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
The second Abelian theorem is rarely mentioned, but see [2] for a simple proof. It is given only for the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
@@
$\lim_{x\nearrow 1}F_{U}(x)=\pm\infty\,$,
with equal signs.
The third Abelian theorem allows only nonnegative real coefficients. It has the advantage over the previous two theorems that it is an equivalence. Now one can go not only from the sum to the functional limit, but also in the reverse. We omit the easy proof.
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist (and are possibly $\infty$) and are equal,
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
@@
Taking the limit for $n\to\infty$ we get that that $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Here the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”. We will use such abbreviated formulations. This result can be easily generalized to linear combinations with more than two terms. Next we consider products of formal power series.
@@
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1$, $n\in\mathbb{N}$, formal power series. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}_{<1}$
@@
$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}F_{U}(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and dividing by $V(1)_{1}$ gives the stated formula. $\Box$
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.19, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Dividing by $W(1)_{1}$ we get the stated formula. $\Box$
We obtain analogs of the previous results in absolute convergence.
@@
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
@@
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}\,$,}
where the inner $\sum$ means sum, the outer $\sum$ means series, and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
For nonnegative real coefficients we can go in the reverse, from the product to the factors.
@@
Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]$. Suppose that $U(x)=V(x)W(x)$, that $U(x)$, $V(x)$ and $W(x)$ are in $H_{1}$, $V(x)W(x)\neq 0$ and that $U(1)_{2}$ exists. Then $V(1)_{2}$ and $W(1)_{2}$ exist, and
@@
Proof. By Theorem 2.18 and the assumption $V(x)W(x)\neq 0$ we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}F_{V}(x)>0$ and the same for $W(x)$. Since $F_{U}=F_{V}F_{W}$ on $\mathbb{C}_{<1}$, we have
$U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}F_{U}(x)$
$=$ $\lim_{x\nearrow 1}F_{V}(x)\cdot\lim_{x\nearrow 1}F_{W}(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
In the second equality we used Theorem 2.18. The third equality holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to F. Mertens.
Theorem 2.26 (Mertens’s)
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
@@
Proof. 1. From the assumption on $c$ it follows that $V(1)_{2}$ exists. From
@@
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
By Proposition 2.6, it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum
@@
Proposition 2.4 shows that $S$ also has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
@@ 2. Indeed, the sum
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By Vandermonde’s identity ([30]),
@@
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of the identity gives
@@
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
beginning with $l$ summands 0. We show that it absolutely converges. Let $T$ be a finite subseries of $S$. Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and
@@
In conclusion, we consider $\xi$-convex light weights.

### v8

@@ 2.5 Sums $U(1)_{1}$ and $U(1)_{2}$ for a generating function $U(x)$
It is the ray in $\mathbb{C}$ emanating from the origin in the direction $\xi$. It is closed under addition. If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$ then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$. If
@@
is a light weight, $n\in\mathbb{N}_{0}$ and $X\subset D_{n}$, then $h(X)\in\xi^{n}\mathbb{R}_{\geq 0}$. The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties. For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
@@
    absolutely converges, we denote its sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
@@
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of Hadamard’s formula for the radius of convergence.
@@
2.6 Three Abelian theorems
We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems. We state them only for the radius of convergence at least $1$. The first theorem is classical, see [1, 2] and [30, Chapter II.7].
@@ 2.6 Three Abelian theorems
$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
The second theorem is seldom mentioned, but see [2] for a simple proof. It is given there only in the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
@@
$\lim_{x\nearrow 1}U(x)=U(1)_{1}=\pm\infty\,.$
The third theorem admits only nonnegative real coefficients. It has two advantages over the previous Abelian theorems: it works for every $U(x)$ in $\mathbb{R}_{\geq 0}[[x]]$ and it is an equivalence, one can go not only from the sum to the limit, but also in the reverse. We omit the easy proof.
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist, are possibly $\infty$ and are equal,
$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
2.7 Conversion of identities in conditional convergence
We believe that the precise statements and proofs of conversions of formal identities to identities at $x=1$, given in this and the next subsection, is our original contribution to the subject.
@@ 2.7 Conversion of identities in conditional convergence
Taking the limit for $n\to\infty$ we get $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
In this proposition, the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates the formulation “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”. We freely use such abbreviations. This result easily generalizes to linear combinations with more terms.
@@
Let $U(x)$, $V_{1}(x)$, $\dots$, $V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1\geq 2$ formal power series. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and if $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
$U(x)=\prod_{i=1}^{n}V_{i}(x)$ for every number $x\in\mathbb{C}_{<1}$
@@
$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}V_{i}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}V_{i}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$, and by dividing by it we get the stated identity. $\Box$
@@
Proof. Formally $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.19, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Dividing by $W(1)_{1}$ we get the stated formula. $\Box$
2.8 Conversion of identities in absolute convergence
We obtain analogs of the results in the previous subsection for absolute convergence.
@@ 2.8 Conversion of identities in absolute convergence
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the stated formula follows from Proposition 2.2. $\Box$
@@
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\times\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
By Proposition 2.5, the series $S$ absolutely converges and has the sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}=\sum_{n\geq 0}u_{n}\,.$}
Here the inner $\sum$ means sum and the outer $\sum$ means series. By Proposition 2.4, the series $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}=V(1)_{2}W(1)_{2}$.
@@ 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
For nonnegative coefficients one can go from the product to the factors.
@@
Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. If $U(x)=V(x)W(x)$, $V(x)W(x)\neq 0$ and $U(1)_{2}$ exists, then $V(1)_{2}$ and $W(1)_{2}$ exist, and
@@
Proof. By Theorem 2.18 and by the assumption $V(x)W(x)\neq 0$, we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}V(x)>0$, and the same applies for $W(x)$. Since $U(x)=V(x)W(x)$ for every $x\in\mathbb{C}_{<1}$, we have
$(+\infty>)\ U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}(V(x)W(x))$
$=$ $\lim_{x\nearrow 1}V(x)\cdot\lim_{x\nearrow 1}W(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
In particular, $V(1)_{2},W(1)_{2}<+\infty$. In the second and fifth equalities we used Theorem 2.18. In the fourth equality we used the arithmetic of functional limits. It holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of this results is due to F. Mertens.
Theorem 2.26 (Mertens)
@@
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
@@
Proof. 1. It follows from the assumption on $c$ that $V(1)_{2}$ exists. From
@@
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,j_{2},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
$S_{m}$ absolutely converges by Proposition 2.6 and has sum $(V(1)_{2}-1)^{m}$. By Propositions 2.2 and 2.4, the series $S$ has sum
@@
Proposition 2.4 shows that $S$ has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
@@ 2. Indeed, the sum
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. Vandermonde’s identity [31] states that
@@
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of Vandermonde’s identity gives
@@
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,$,}
beginning with $l$ zero summands. We show that it absolutely converges. Let $T$ be a finite subseries of $S$. Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=|v|+|v|\sum_{m\geq 1}c^{m}=\frac{|v|}{1-c}\,.$}
Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
@@
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and that $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and if
@@
2.9 Generating functions of walks for $\xi$-convex weights
We generalize the result in Section 1.2 that $h(D_{n})=1$ for every $n$.